# Quenched coalescent for diploid population models with selfing and overlapping generations

### v1

@@ 5.4 Mixed demographies
Combining Equations (5.4), (5.5), and (5.6) shows that $Q_{n}^{(2)}$ corresponds exactly to the measure $\Xi^{(2)}=\delta_{\left(\frac{\psi}{2},\frac{\psi}{2},0,\ldots\right)}$. The result then follows from Corollary 5.6. ∎
@@
Possible references for what follows are BK10a ; BK10b ; Led (16); Mit (83).
To prove Theorem 4.7 we require a characterization of weak convergence in distribution of random measures. Following (Kal, 17, Theorem 4.19, pp. 111, 126), we provide the following definition of weak convergence in distribution for random probability measures.
@@
Let $\mu$ and $(\mu_{N})_{N\in\mathbb{N}}$ denote random probability measures on a Polish space $S$. Then $\mu_{N}$ converges weakly in distribution to $\mu$ if for any continuous $f$ of bounded support that
$\mu_{N}(f)\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{d}}}}\,\mu(f).$
When this convergence holds we write $\mu_{N}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd}}}}\,\mu$.
We provide in this section sufficient criteria to characterize weak convergence in distribution when the Polish space $S$ in question is precisely $\mathcal{D}\left(\mathbb{R}_{+},E\right)$ with the $J_{1}$ topology, where $E$ is a locally compact Polish space. The basic strategy is to construct a suitable dense family in $C_{b}\left(\mathcal{D}\left(\mathbb{R}_{+},E\right)\right)$ and then to devise criteria by which evaluations our random measures against this suitable class of test functions converge.
@@ 6.1 A dense family in $C_{b}\left(\mathcal{D}\left(\mathbb{R}_{+},E\right)\right)$
Suppose that $E$ is a locally compact Polish space. Then $\mathcal{D}\left(\mathbb{R}_{+},E\right)$ is a Polish space under the $J_{1}$ topology (EK, 09, Theorem 5.6, p.121). We shall describe a dense subset of $C_{b}\left(\mathcal{D}\left(\mathbb{R}_{+},E\right)\right)$ in Lemma 6.4 below when $C_{b}\left(\mathcal{D}\left(\mathbb{R}_{+},E\right)\right)$ is given the topology of uniform convergence on compacta. To this end, we first describe an explicit continuous function on $\mathcal{D}\left(\mathbb{R}_{+},E\right)$.
@@
where each $I_{ij}$ belongs to $\mathcal{I}$ and $k_{i},n\in\mathbb{Z}_{+}$. An extension of this class of continuous, bounded integral functionals will, by an application of Stone-Weierstrass Ok (25)[Chapter 7, p. 11], be dense in $C_{b}(\mathcal{D}\left(\mathbb{R}_{+},E\right))$.
@@
Fix any compact set $K\subset\mathcal{D}\left(\mathbb{R}_{+},E\right)$ and any $f$ in $C_{b}(\mathcal{D}\left(\mathbb{R}_{+},E\right))$. The restriction of elements of $\bar{\mathcal{I}}$ to $K$ yields an algebra $\bar{\mathcal{I}}(K)$ of continuous functions from $K$ to $\mathbb{R}$ that contains the constants and separates points. By the Stone-Weierstrass theorem for compact Hausdorff spaces Ok (25)[Chapter 7, p. 11], the uniform closure of $\bar{\mathcal{I}}(K)$ is all of $C(K)$. This gives the claim.
@@ 6.2 A characterization of weak convergence in distribution
Let $\mu_{N}$ denote a sequence of random measures in $\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},E\right)\right)$ for a locally compact, Polish space $E$. Then $\mu_{N}$ converges weakly in distribution to $\mu$ if and only if for any finite collection $\{\varphi_{i}\}_{i=1}^{k}$ of elements in $C_{c}\left(\mathbb{R}_{+}\times E\right)$ that
@@
From Definition 6.1 it suffices to show that for any $f$ in $C_{b}\left(\mathcal{D}\left(\mathbb{R}_{+},E\right)\right)$ that
$\mu_{N}(f)\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{d}}}}\,\mu(f).$ (6.2)
@@
The characterization of weak convergence in distributions of random measures on $\mathcal{D}\left(\mathbb{R}_{+},E\right)$ for locally compact and Polish $E$ given by Theorem 6.5 allows one to give classical-like characterizations of this form of weak convergence, as typified by the following following.
@@
Let $\mu$, $(\mu_{N})_{N\in\mathbb{N}}$ denote random probability measures over the space $\mathcal{D}\left(\mathbb{R}_{+},E\right)$. The sequence $\mu_{N}$ converges weakly in finite-dimensional distribution to $\mu$ if for any finite collection of times $\left(t_{i}\right)_{1\leq i\leq k}$ the pushforward by the evaluation map
@@
converges weakly in distribution as random measures. That is,
$\left(\pi_{\vec{t}}\right)_{*}\mu_{N}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd}}}}\,\left(\pi_{\vec{t}}\right)_{*}\mu.$
We know define an analogue of the compact containment condition for random measures on Skorokhod space (EK, 09, p. 129).
@@
Let $\mu_{N}$ denote a sequence of random measures on $\mathcal{D}\left(\mathbb{R}_{+},E\right)$. Then $\mu_{N}$ satisfies the compact containment condition if, for every $\varepsilon>0$ and $T>0$ that there is a compact set $K(\varepsilon,T)\subset E$ such that
$\mathbb{P}\left(\limsup_{N\to\infty}\mu_{N}\left(x([0,T])\subset K(\varepsilon,T)\right)\geq 1-\varepsilon\right)=1.$
We obtain the following characterization of weak convergence of random measures in $\mathcal{D}\left(\mathbb{R}_{+},E\right)$.
@@
Suppose $\mu$ is a random measure on $\mathcal{D}\left(\mathbb{R}_{+},E\right)$, where $E$ is a locally compact Polish space. Suppose now that $\mu_{N}$ is a sequence of random measures on $\mathcal{D}\left(\mathbb{R}_{+},E\right)$ such that
  • for any finite collection of times $\left(t_{i}\right)_{1\leq i\leq k}$ we have that $\left(\pi_{\vec{t}}\right)_{*}\mu_{N}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd}}}}\,\left(\pi_{\vec{t}}\right)_{t}\mu$,
  • and $\mu_{N}$ satisfies the compact containment condition for random measures from Definition 6.7,
@@
Since $\varphi_{i}$ is continuous on $0,T]\times K(\eta,T)$ for any compact $K\subset E$, it is uniformly continuous there. Fix now $\delta>0$. Choose a fine partition $0=t_{0}<t_{1}<\ldots<t_{r}=T$ so that $\min_{i}\{t_{i+1}-t_{i}\}<\delta$. By uniform continuity, for any $\varepsilon>0$ we can choose $\delta$ so that
@@
for any $|s-t|<\delta$.
For any path $x$ for which $x(s)\in K(\eta,T)$ for all $0\leq s\leq T$ we have that
$\left|\int_{0}^{T}\varphi_{i}\left(s,x(s)\right)-\sum_{j=0}^{r-1}\left(t_{j+1}-t_{j}\right)\varphi_{i}(t_{j},x(t_{j}))\right|\leq T\varepsilon.$ (6.5)
For all other $x$ we have that
@@
Combining (6.5), (6.6) yields
@@
By Markov’s inequality and (6.4) we therefore have, for any fixed $\theta>0$, that
@@
As $\eta,\rho,\varepsilon,\theta$ are arbitrary, we therefore have that
$\mu_{N}\left(\left(\left|\int_{0}^{T}\varphi_{i}\left(s,x(s)\right)-\sum_{j=0}^{r-1}\left(t_{j+1}-t_{j}\right)\varphi_{i}(t_{j},x(t_{j}))\right|\right)_{1\leq i\leq k}\in\cdot\right)\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd}}}}\,\delta_{0}$
as $N$ goes to infinity.
As $\left(\pi_{\vec{t}}\right)_{*}\mu_{N}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd}}}}\,\left(\pi_{\vec{t}}\right)_{*}\mu$, the vector of Riemann sums satisfies
$\left(\pi_{\vec{t}}\right)_{*}\mu_{N}\left(\left(\sum_{j=0}^{r-1}\left(t_{j+1}-t_{j}\right)\varphi_{i}(t_{j},x(t_{j}))\right)_{1\leq i\leq k}\in\cdot\right)\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd}}}}\,\left(\pi_{\vec{t}}\right)_{*}\mu\left(\left(\sum_{j=0}^{r-1}\left(t_{j+1}-t_{j}\right)\varphi_{i}(t_{j},x(t_{j}))\right)_{1\leq i\leq k}\in\cdot\right).$
This gives the claim.
@@
Corollary 6.9 allows us to strengthen some already existing results, including (ABFW, 25, Theorem 3.8). Theorem 3.8 demonstrates quenched convergence in finite-dimensional distribution to a time-inhomogeneous coalescent. By the corollary this is enough to show that the quenched convergence in laws follows.
@@ 7.1 Construction of the $Q$-$\lambda$ graph
Fix $\lambda\in\mathbb{R}_{+}$ and a sequence of consistent matrices $\left(Q_{n}\right)_{n\in\mathbb{N}}$, and write $Q_{n}=\left(q_{\xi\eta}^{n}\right)_{\xi,\eta\in\mathcal{E}_{n}}$. We can assume they are obtained from Assumptions 3.4 and 3.6 respectively. We shall define a stochastic process, called a a $Q$-$\lambda$ ancestral graph, with state space
$\Theta:=\mathbb{Z}_{+}\times\mathbb{Z}_{+}\times\mathcal{E},$
where $\mathcal{E}=\cup_{n\geq 2}\mathcal{E}_{n}$. For each $(l,m,\xi)\in\Theta$, the integer $l$ will track the current number of lineages, $m$ the label of the particle that most recently fragmented, and $\xi$ the structure of the most recent coalescence event.
@@
A $Q$-$\lambda$ ancestral graph is a continuous-time Markov process with sample paths in $\mathcal{D}\left(\mathbb{R}_{+},\Theta\right)$ and infinitesimal generator $\mathscr{L}$ acting on $f\in C_{b}(\Theta)$ by
$\mathscr{L}f(l,m,\xi)=\sum_{j=1}^{l}\lambda\left[f(l+1,j,\xi)-f(l,m,\xi)\right]+\sum_{\eta\in\mathcal{E}_{l}\setminus\{\xi_{0}^{l}\}}q_{\xi_{0}^{l}\eta}^{l}\left[f(|\eta|,m,\eta)-f(l,m,\xi)\right]$ (7.1)
@@
Existence of $G_{Q,\lambda}^{n}$ holds because fragmentation rate is linear in the first coordinate and does not depend on the the other two coordinates. This can be verified by, for instance, (EK, 09, Proposition 2.9 in Chap. 4).
We describe now $n$ independent random walks $\{x_{i}\}_{1\leq i\leq n}$ on the ancestral graph $G_{Q,\lambda}^{n}$, which we view as $\mathbb{Z}_{+}$-valued processes. The joint process $\left(G_{Q,\lambda}^{n},x_{1},\ldots,x_{n}\right)$ may be described by a generator $\tilde{\mathscr{L}}$ defined on $C_{b}\left(\Theta\times\mathbb{Z}_{+}^{n}\right)$. Define a vector $\sigma:\mathbb{Z}_{+}^{n}\times\mathbb{Z}_{+}\to\{0,1\}^{n}$ by
@@
The $n$ $\mathbb{Z}_{+}$-valued random walks $\{x_{i}\}_{1\leq i\leq n}$ on the $Q$-$\lambda$ ancestral graph $G_{Q,\lambda}^{n}$ have their joint distribution with $G_{Q,\lambda}^{n}$ described by the generator $\tilde{\mathscr{L}}$. They satisfy that

### v2

@@ 5.4 Mixed demographies
Combining Equations (5.9), (5.10), and (5.11) shows that $Q_{n}^{(2)}$ corresponds exactly to the measure $\Xi^{(2)}=2\delta_{\left(\frac{\psi}{2},\frac{\psi}{2},0,\ldots\right)}$ on $\Delta\setminus\{\bf{0}\}$. ∎
@@
To prove Theorem 4.7 we shall develop a characterization of weak convergence in distribution of random probability measures in Theorem 6.5. We are not aware of any previous recognition of this characterization in the literature, so Theorem 6.5 may have general interest.
More precisely, we provide in this section sufficient criteria to characterize weak convergence in distribution when the Polish space $S$ in question is precisely $\mathcal{D}\left(\mathbb{R}_{+},E\right)$ with the $J_{1}$ topology, where $E$ is a locally compact Polish space (such as $\mathcal{E}_{n}$ or $\mathcal{E}$). The basic strategy is to construct a suitable dense family in $C_{b}\left(\mathcal{D}\left(\mathbb{R}_{+},E\right)\right)$ and then to devise criteria by which evaluations our random measures against this suitable class of test functions converge.
Let $\mathcal{S}$ be a Polish space, such as $\mathcal{D}\left(\mathbb{R}_{+},E\right)$. Under the weak topology on $\mathcal{M}_{1}\left(\mathcal{S}\right)$, convergence in distribution (denoted by $\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{d}}}$}\, in this paper) of a sequence $(\mu_{N})_{N\in\mathbb{N}}$ of random elements in $\mathcal{M}_{1}\left(\mathcal{S}\right)$ is also called “weak convergence in distribution” and is denoted by $\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd\;}}}$}\, in (Kal, 17, Chapter 4). That is,
@@
Let $\mu$ and $(\mu_{N})_{N\in\mathbb{N}}$ be random probability measures on a Polish space $\mathcal{S}$. We say that $\mu_{N}$ converges weakly in law to $\mu$ and write
$\mu_{N}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd\;}}}}\,\mu\quad\text{in}\quad\mathcal{M}_{1}(\mathcal{S})$
if $\mu_{N}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{d}}}}\,\mu$ when $\mathcal{M}_{1}(S)$ is equipped with the weak topology.
By (Kal, 17, Theorem 4.19)
$\mu_{N}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd\;}}}}\,\mu\text{\quad if and only if\quad}\mu_{N}(f)\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{d}}}}\,\mu(f)$ (6.1)
for any continuous and bounded function of bounded support $f$. We use the latter equivalent characterization throughout for the Polish space $\mathcal{S}=\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)$.
@@ 6.1 A dense family in $C_{b}\left(\mathcal{D}\left(\mathbb{R}_{+},E\right)\right)$
Suppose that $E$ is a locally compact Polish space. Then $\mathcal{D}\left(\mathbb{R}_{+},E\right)$ is a Polish space with the $J_{1}$ topology (EK, 09, Theorem 5.6, p.121), under a bounded metric; see (EK, 09, Chapter 3, eqn. (5.2)).
We shall describe a dense subset of $C_{b}\left(\mathcal{D}\left(\mathbb{R}_{+},E\right)\right)$ in Lemma 6.4 below when $C_{b}\left(\mathcal{D}\left(\mathbb{R}_{+},E\right)\right)$ is given the topology of uniform convergence on compacta. To this end, we first describe an explicit continuous function on $\mathcal{D}\left(\mathbb{R}_{+},E\right)$.
@@
where each $I_{ij}$ belongs to $\mathcal{I}$ and $k_{i},n\in\mathbb{Z}_{+}$. An extension of this class of continuous, bounded integral functionals will, by an application of Stone-Weierstrass (Ok, 25, Chapter 7, p. 11), be dense in $C_{b}(\mathcal{D}\left(\mathbb{R}_{+},E\right))$.
@@
Fix any compact set $K\subset\mathcal{D}\left(\mathbb{R}_{+},E\right)$ and any $f\in C_{b}(\mathcal{D}\left(\mathbb{R}_{+},E\right))$. The restriction of elements of $\bar{\mathcal{I}}$ to $K$ yields an algebra $\bar{\mathcal{I}}(K)$ of continuous functions from $K$ to $\mathbb{R}$ that contains the constants and separates points. By the Stone-Weierstrass theorem for compact Hausdorff spaces (Ok, 25, Chapter 7, p. 11), the uniform closure of $\bar{\mathcal{I}}(K)$ is all of $C(K)$. This gives the claim.
@@ 6.2 A characterization of weak convergence in distribution
Let $\left(\mu_{N}\right)_{N\in\mathbb{N}}$ and $\mu$ be random measures taking value in $\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},E\right)\right)$, where $E$ is a locally compact Polish space. Then $\mu_{N}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd\;}}}}\,\mu$ if and only if for any finite collection $\{\varphi_{i}\}_{i=1}^{k}$ of elements in $C_{c}\left(\mathbb{R}_{+}\times E\right)$ that
@@
We first prove sufficiency of (6.2). From the ”if part” of (6.1), it suffices to show that
$\mu_{N}(f)\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{d}}}}\,\mu(f)\quad\text{for }f\in C_{b}\left(\mathcal{D}\left(\mathbb{R}_{+},E\right)\right).$ (6.3)
@@
The characterization of weak convergence in distributions of random measures on $\mathcal{D}\left(\mathbb{R}_{+},E\right)$ for locally compact and Polish $E$ given by Theorem 6.5 allows one to give classical-like characterizations of such convergence, as typified by Theorem 6.8 below.
@@
Let $\left(\mu_{N}\right)_{N\in\mathbb{N}}$ and $\mu$ be random measures taking value in $\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},E\right)\right)$, where $E$ is a locally compact Polish space. We say that the sequence $\mu_{N}$ converges weakly in finite-dimensional distribution to $\mu$ if
$\left(\pi_{\vec{t}}\right)_{*}\mu_{N}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd\;}}}}\,\left(\pi_{\vec{t}}\right)_{*}\mu\quad\text{in}\quad\mathcal{M}_{1}(\mathbb{R}^{k})$
for any finite collection of times $\left(t_{i}\right)_{1\leq i\leq k}$, for any $k\geq 1$, where $\pi_{\vec{t}}$ is the evaluation map
@@
and $\left(\pi_{\vec{t}}\right)_{*}\mu$ is the pushforward of $\mu$ by $\pi_{\vec{t}}$.
We now define an analogue of the compact containment condition for random measures on Skorokhod space (EK, 09, p. 129).
@@
Let $\{\mu_{N}\}$ denote a sequence of random measures on $\mathcal{D}\left(\mathbb{R}_{+},E\right)$. Then $\{\mu_{N}\}$ satisfies the compact containment condition if, for every $\varepsilon>0$ and $T>0$ that there is a compact set $K(\varepsilon,T)\subset E$ such that
$\mathbb{P}\left(\liminf_{N\to\infty}\mu_{N}\left(x([0,T])\subset K(\varepsilon,T)\right)\geq 1-\varepsilon\right)=1.$
We obtain the following characterization of weak convergence in law of random measures in $\mathcal{D}\left(\mathbb{R}_{+},E\right)$.
@@
Let $E$ be a locally compact Polish space. Let $\mu$ and $\{\mu_{N}\}_{N\in\mathbb{N}}$ be random measures taking value in $\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},E\right)\right)$. Suppose the followings hold:
  • (i)
    $\{\mu_{N}\}$ converges weakly in finite-dimensional distribution to $\mu$ as in Definition 6.6, and
  • (ii)
    $\{\mu_{N}\}$ satisfies the compact containment condition as in Definition 6.7.
@@
To this end, let $F_{i}$ be defined by
$F_{i}:=\int_{0}^{\infty}\varphi_{i}(s,x(s))ds.$
Fix $G$ in $C_{b}(\mathbb{R}^{k})$. We need to show that
$\mathbb{E}\left[G\left(\mu_{N}(F_{1}),\ldots\mu_{N}(F_{k})\right)\right]\to\mathbb{E}\left[G\left(\mu(F_{1}),\ldots\mu(F_{k})\right)\right].$
We will proceed by approximating $F_{i}$ by Riemann sums.
@@
Since $\varphi_{i}$ is continuous on $0,T]\times K(\eta,T)$ for any compact $K\subset E$, it is uniformly continuous there. Fix $\varepsilon>0$. By uniform continuity, for any $\varepsilon>0$ we can choose $\delta>0$ so that
@@
for every $i=1,\ldots,k$ and any $|s-t|<\delta$. Choose a fine partition $0=t_{0}<t_{1}<\ldots<t_{r}=T$ so that $\min_{i}\{t_{i+1}-t_{i}\}<\delta$ and define the discrete Riemann sum $R_{i}^{\varepsilon}$ by
$R_{i}^{\varepsilon}:=\sum_{j=0}^{r-1}\left(t_{j+1}-t_{j}\right)\varphi_{i}(t_{j},x(t_{j})).$
For ease of notation we write $\mu_{N}(F)$ for the tuple
$\left(\mu_{N}(F_{1}),\ldots,\mu_{N}(F_{k})\right)$
in $\mathbb{R}^{k}$, and similarly for $\mu_{N}(R^{\varepsilon})$, $\mu(F)$, and $\mu(R^{\varepsilon})$. By the triangle inequality we have that
$\left|\mathbb{E}\left[G(\mu_{N}(F))-G(\mu(F))\right]\right|$ (6.6)
$\leq$ $\mathbb{E}\left[\left|G(\mu_{N}(F))-G(\mu_{N}(R^{\varepsilon}))\right|\right]+\left|\mathbb{E}\left[G(\mu_{N}(R^{\varepsilon}))-G(\mu(R^{\varepsilon}))\right]\right|+\mathbb{E}\left[\left|G(\mu(R^{\varepsilon}))-G(\mu(F))\right|\right].$ (6.7)
We first show that the second summand in (6.6) can be made arbitrarily small. Recall that $\{t_{i}\}_{i=0}^{r}$ are the points of our mesh for the Riemann sums $R^{\varepsilon}$ with $t_{0}=0$ and $t_{r}=T$. Define the map $H_{i}:E^{r}\to\mathbb{R}$ by
$H_{i}(e_{0},\ldots,e_{r}):=\sum_{j=0}^{r-1}(t_{j+1}-t_{j})\varphi_{i}(t_{j},e_{j}).$
Then $R_{i}^{\varepsilon}=H_{i}\circ\pi_{\vec{t}}$. As $\mu_{N}$ converges weakly in finite-dimensional distribution by assumption (i), $\mu_{N}(R^{\varepsilon})$ converges in distribution to $\mu(R^{\varepsilon})$. Therefore the second summand in (6.6) can be made arbitrarily small..
@@
$|G(u)-G(v)|<\delta\text{\, when \,}\max_{1\leq i\leq k}|u_{i}-v_{i}|<\gamma.$
By uniform continuity of $G$ on $M,M]^{k}$ we therefore have that
@@
Taking expectations over (6.8) yields
@@
As we can take $\delta$ arbitrarily small, it will suffice to show that the probability term can be made arbitrarily small.
To show the probability term can be made arbitrarily small, we utilize the compact containment condition. Fix $\eta,\rho>0$. By assumption (ii), there exists a compact set $K(\eta,T)$ such that, for the event $A_{N}:=\{\mu_{N}(x([0,T])\subset K(\eta,T))\geq 1-\eta\}$
$\mathbb{P}\left(A_{N}\right)\geq 1-\rho$
for all sufficiently large $N$. On $A_{N}$ we therefore have
[…]
As $\eta$ and $\varepsilon$ we can take arbitrarily small, we can take them so that
[…]
for all $i=1,\ldots,k.$ Therefore
$\mathbb{P}\left(\max_{i}\left|\mu_{N}(F_{i})-\mu_{N}(R_{i}^{\varepsilon})\right|\geq\gamma\right)\leq\mathbb{P}\left(A_{N}^{c}\right)\leq\rho$ (6.10)
for all sufficiently large $N$. Combining (6.10) with (6.9) gives
[…]
Taking $\delta$ and $\rho$ arbitrarily small then makes the first summand of (6.6) arbitrarily small.
The third summand can be made arbitrarily small by the same argument as for the first summand, giving the result.
@@
Corollary 6.9 allows us to strengthen some already existing results, including (ABFW, 25, Theorem 3.8) which demonstrates weak convergence in finite-dimensional distribution. By Corollary 6.9, weak convergence in law therefore follows from compactness of $\mathcal{E}_{n}$.
@@ 7.1 Construction of the $Q$-$\lambda$ graph
Fix $\lambda\in\mathbb{R}_{+}$ and a sequence of consistent matrices $\left(Q_{n}\right)_{n\in\mathbb{N}}$, and write $Q_{n}=\left(q_{\xi\eta}^{n}\right)_{\xi,\eta\in\mathcal{E}_{n}}$. We assume they are obtained from Assumptions 3.5 and 3.7 respectively. We define a stochastic process, called a a $Q$-$\lambda$ ancestral graph, with state space
$\Omega:=\mathbb{Z}_{+}\times\mathbb{Z}_{+}\times\mathcal{E},$ (7.1)
where $\mathcal{E}=\cup_{n\geq 2}\mathcal{E}_{n}$. For each $(l,m,\xi)\in\Omega$, the integer $l$ will track the current number of lineages, $m$ the label of the particle that most recently fragmented, and $\xi$ the structure of the most recent coalescence event.
@@
A $Q$-$\lambda$ ancestral graph is a continuous-time Markov process with sample paths in $\mathcal{D}\left(\mathbb{R}_{+},\Omega\right)$ and infinitesimal generator $\mathfrak{L}$ acting on $f\in C_{b}(\Omega)$ by
$\mathfrak{L}f(l,m,\xi)=\sum_{j=1}^{l}\lambda\left[f(l+1,j,\xi)-f(l,m,\xi)\right]+\sum_{\eta\in\mathcal{E}_{l}\setminus\{\xi_{0}^{l}\}}q_{\xi_{0}^{l}\eta}^{l}\left[f(|\eta|,m,\eta)-f(l,m,\xi)\right]$ (7.2)
@@
Existence of $G_{Q,\lambda}^{n}$ holds because the fragmentation rate is linear in the first coordinate and does not depend on the the other two coordinates. This can be verified by, for instance, (EK, 09, Proposition 2.9 in Chap. 4).
Next we describe $n$ independent random walks $\{x_{i}\}_{1\leq i\leq n}$ on the ancestral graph $G_{Q,\lambda}^{n}$, which we view as $\mathbb{Z}_{+}$-valued processes. The joint process $\left(G_{Q,\lambda}^{n},x_{1},\ldots,x_{n}\right)$ may be described by a generator $\tilde{\mathfrak{L}}$ defined on $C_{b}\left(\Omega\times\mathbb{Z}_{+}^{n}\right)$. Define a vector $\sigma:\mathbb{Z}_{+}^{n}\times\mathbb{Z}_{+}\to\{0,1\}^{n}$ by
@@
The $n$ $\mathbb{Z}_{+}$-valued random walks $\{x_{i}\}_{1\leq i\leq n}$ on the $Q$-$\lambda$ ancestral graph $G_{Q,\lambda}^{n}$ have their joint distribution with $G_{Q,\lambda}^{n}$ described by the generator $\tilde{\mathfrak{L}}$. They satisfy that